OpenStudy (anonymous):

For each quadratic function, find the vertex of its graph. y= -x^2 + 6x - 4

OpenStudy (anonymous):

vertex is at \(-\frac{b}{2a}\) for the x coordinate

OpenStudy (anonymous):

in your case you have \(-\frac{6}{2\times (-1)}=3\) to find the second coordinate, replace x by 3 to see what you get

OpenStudy (anonymous):

So 24